OpenStudy (anonymous):

Find the lim of tan(2x)/(3x+sinx) as x approaches 0

OpenStudy (noelgreco):

I don't see any quick way without l'Hopital.

OpenStudy (anonymous):

l'Hopitals has taken me back to the start

OpenStudy (anonymous):

If l'Hopital doesn't work the first time but gives you the condition for it again, you can repeat it. You might have better luck the second time around.

OpenStudy (anonymous):

1/4

OpenStudy (anonymous):

One way is to use l'Hopital, but a faster way is to use the linear approximations for sine and tangent. The answer is 1/2.

OpenStudy (anonymous):

THE DIFFERENTIAL IS SEC^2(X)/3+COS(x).numerator is 1 and denominator is 4 answer is 1/4 how did you get 1/2??????

OpenStudy (anonymous):

pasta! skywalker94 is right! Using l'Hopital, we'll get 1/2. \[\lim_{x \rightarrow 0}(\tan2x/(3x+sinx))=\lim_{x \rightarrow 0}[2(1+\tan^{2}2x)/(3+cosx)]=1/2\]

OpenStudy (anonymous):

use expansion method it is the fastest.... tanx = x + x^3/3............sinx = x - x^3/3!

OpenStudy (anonymous):

and from that the ans is 1/2 Math Physics and skywalker ur right..!!!!